OpenStudy (anonymous):

Can anyone help me with 2 questions?

OpenStudy (anonymous):

I might be able to

OpenStudy (anonymous):

depends, what are the questions?

OpenStudy (anonymous):

k thanks i'll put the first question

OpenStudy (anonymous):

Elizabeth had a book delivered to her home. The total cost included the price of the book, 7% sales tax on the price of the book, and a $4.50 shipping charge. Elizabeth paid a total of $31.25. What was the pre-tax price of the book?

OpenStudy (anonymous):

i am on unit 9 lesson 11 : Multistep ratio and percent

OpenStudy (anonymous):

i am going to put the other question

OpenStudy (anonymous):

Choose exactly two answers that are correct. David purchased an airline ticket. Included in the total price were $300 for the ticket, 7.5% tax on the price of the ticket, and fees of $27. Which statements are true? Choose exactly two answers that are correct. A. The equation that represents the total price, t, is t + 0.075t + 27 = 300. B. The equation that represents the total price, t, is 300 + 0.075(300) + 27 = t. C. The total cost of the airline ticket is $349.50. D. The total cost of the airline ticket is $253.95.

OpenStudy (anonymous):

This is for question 1: 7% = 0.07 Price of the book = x x + 0.07x + 4.5 = 31.25 1.07x + 4.5 = 31.25 1.07x = 26.75 x = 25 Pre Tax Price = $25

OpenStudy (purple_pink):

k12?

OpenStudy (anonymous):

ya lol

OpenStudy (anonymous):

@The_One_And_Only i thought that was the answer but i wasn't sure

OpenStudy (anonymous):

do you have any idea on the second one

OpenStudy (anonymous):

Im working on it

OpenStudy (anonymous):

k thanks :D

OpenStudy (anonymous):

I think the answer is C and B (you can double check my answer but I think it's right)

OpenStudy (anonymous):

i thought it was a and c but let me double check

OpenStudy (anonymous):

No, because in the question it said INCLUDED in the total price was 300, not the total price WAS 300

OpenStudy (anonymous):

OH NOW I SEE your right!

OpenStudy (anonymous):

let me submit my test and see though

OpenStudy (anonymous):

i got 100% :D

OpenStudy (anonymous):

Awesome!

OpenStudy (anonymous):

thank you for your help!

OpenStudy (anonymous):

No problem, well 2 problems but, your welcome! :D

OpenStudy (anonymous):

i fanned you

OpenStudy (anonymous):

thanks, I did to :)

OpenStudy (anonymous):

bye:D

OpenStudy (anonymous):

bye!